OpenStudy (anonymous):

Write an equation for the line that passes through (0, 9) and whose slope is -3. y = 3x – 9 y = 9x + 3 y = -3x + 9 y = -3x – 9

OpenStudy (ghazi):

y=mx+c .....y=-3x+c now to find c put x=0 and y=9..you'll have c= 9 and your equation is y=3x+9

OpenStudy (ghazi):

sorry y=-3x+9

OpenStudy (anonymous):

yay i also got that thanks man <3

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

Write an equation that passes through (0, -2) and whose slope is ½. y = 1/2x + 2 y = -1/2x – 2 y = 1/2x – 2 y = -2x + 1/2 for thi i got the 3d one is that correct

OpenStudy (ghazi):

C

OpenStudy (anonymous):

wooo thanks so happy i got both thissite really helps

OpenStudy (anonymous):

can you do this btw Find the solution set for 2x – 2y = 8, given the replacement set {(-2, 2), (-1, -5), (1, -3), (2, -2)}. {(1, -5), (1, -3), (2, -2)} {(-2, 2), (-1, -5), (1, -3)} {(-2, -6), (-1, -5), (1, -3), (2, 2)} {(-1, -5), (1, -3)}